If A(-1,4),B(2,3),C(8,1) then which point is between the other two points ???

Grade:11

1 Answers

Ritesh Khatri
6 years ago
By using distance formula
AB = 10^(1/2)
BC = 2*(10^(1/2))
AC = 3*(10^(1/2))
 
By observation AB + BC = AC
Therefore B lie between A and C